Hello … please explain
 
For steel the breaking stress is 8 * 106 N/m2 and the density is 8 * 103 Kg/m3.What is the maximum length of a steel wire which can be suspended without breaking under its own weight? ( 9.8 m/s^2)

Grade:9

1 Answers

Vikas TU
7 years ago
Stress = force/Area
8*10^6 = 9.8*(Density*V)/Area
0.816 x 10^6 = 8*10^3*Area*l/Area
l = 0.102 x 10^3 m